OpenStudy (anonymous):

How do I find the domain of (f*g)(x) for each f(x) and g(x): f(x)=√x-2 g(x)= (1) / (4x) and get x≤ 1/8?

OpenStudy (anonymous):

\[f(x)=\sqrt{x-2}\] since you cannot take the square root of a negative number find the domain via \[x-2\geq 0\] or \[x\geq 2\]

OpenStudy (anonymous):

is \[g(x)=\frac{1}{4x}\]?

OpenStudy (anonymous):

also, is it ] \[f\times g(x)\] or \[f\circ g(x)\]?

OpenStudy (anonymous):

the second one. f∘g(x)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Basically its asking you for \[f(g(x))\] .

OpenStudy (anonymous):

ooh ok that makes more sense

OpenStudy (anonymous):

\[\large f\circ g(x)=f(g(x))=f(\frac{1}{4x})=\sqrt{\frac{1}{4x}-2}\]

OpenStudy (anonymous):

since you cannot take the square root of a negative number, your job is to solve \[\frac{1}{4x}-2\geq 0\]

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good you should get \[0<x\leq \frac{1}{8}\]

OpenStudy (anonymous):

yup thats what i got! thanks!

OpenStudy (anonymous):

yw